MEGHÍVÓ

2004. február 9-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

ADRIAN MATHIAS (Cambridge, Anglia)

TWO INTUITIONS, TWO FOUNDATIONS

című előadása.

Of the responses to the collapse in 1900 of Frege’s theory of classes, two are going strong a century later: type theory and set theory. This talk will compare the strengths and weaknesses of their respective claims to provide a foundation for mathematics.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2004. február 9-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

ADRIAN MATHIAS (Cambridge, Anglia)

TWO INTUITIONS, TWO FOUNDATIONS

című előadása.

Of the responses to the collapse in 1900 of Frege’s theory of classes, two are going strong a century later: type theory and set theory. This talk will compare the strengths and weaknesses of their respective claims to provide a foundation for mathematics.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!